HUBBLE SPACE TELESCOPE DAILY REPORT #2152 PERIOD COVERED: 0000Z (UTC) 06/24/98 - 0000Z (UTC) 06/25/98 Daily Status Report as of 176/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed STIS/CCD 7926 (CCD Dark and Bias Monitor -- Continued) The Space Telescope Imaging Spectrograph (CCD) was used to make several dark and bias calibration observations. This is to measure CCD dark current and bias rates for the two supported amplifier settings {gain = 1 and gain = 4} in order to gather data for making superdarks and superbiases. These observations will also track the growth of hot pixels on the CCD in general and within the various target acquisition apertures specifically. The observations were executed as scheduled, and no anomalies were noted. 1.2 Completed STIS/CCD 7645 (STIS FUV-MAMA Cycle 7 Flats) The Space Telescope Imaging Spectrograph (CCD) was used to make far-ultraviolet flat calibration images. This program will obtain FUV-MAMA flat-field observations with the Kr lamp for the construction of on-orbit D-flats for G140M. A dark calibration image was also taken. The observations were completed as planned, and no anomalies were reported. 1.3 Completed Three Sets of STIS/CCD 7908 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal) The Space Telescope Imaging Spectrograph (CCD) was used to make calibration observations as part of the POMS Test Proposal. This was a STIS non-scripted parallel proposal. The observations were completed as planned, and no problems were reported. 1.4 Completed NIC/3 7841 (HST/NICMOS Observations of the Nearest Embedded Cluster: the L1688 {Rho Oph} Protostellar Region) The Near Infrared Camera (Camera Three) was used to make observations protostellar region of the nearest young cluster L1688 (positions 1, 2, 3 and 4) ( a.k.a., Rho Oph). This cluster is unique in its proximity to Earth, only 125 pc away; and in its high proportion of optically invisible protostars. The observations were executed as scheduled, and no anomalies were noted. 1.5 Completed WF/PC-2 8049 (Long WF/PC-2 Decontaminations during 2nd NIC/3 Campaign) The WF/PC-2 was used to make long decontamination procedures as part of the second NIC/3 campaign. Each decon is followed by five 30-minute darks. During this time, no external science WF/PC-2 observations were executed. The decontaminations were executed as scheduled, and no anomalies were noted. 1.6 Completed NIC/1/2/3 7830 (IR Spectra for Known-Mass M Dwarfs) The Near Infrared Camera (Camera One, Two and Three) was used to make infrared spectrographic observations of M-Class dwarf star J1228-1547. Over the last decade there has been considerable debate on two major issues in the low-mass star literature: {1} the effective temperature scale for M dwarfs and its relation to mass and luminosity and {2} the conversion from an absolute magnitude in some passband to a bolometric magnitude. The observations were executed as scheduled, and no anomalies were noted. 1.7 Completed NIC/1/2/3 7919 (NICMOS Camera 3 Campaign Snapshots) The Near Infrared Camera (Camera Three) was used to make images of galaxy NGC-976 as part of the NICMOS Camera 3 campaign snapshots. The proposal is to obtain wide field snapshot images of nearby galaxies in Paschen Alpha which is inaccessible from the ground. The observations were executed as planned, and no anomalies were reported. 1.8 Completed NIC/1/2/3 7967 (HST Observations of a Galaxy at Z=5.34) The Near Infrared Camera (Cameras One, Two and Three) was used to make observations of galaxy 0140+326RD which is located at a distance of Z = 5.34. This galaxy is the only spectroscopically confirmed object at z>5. The observations were executed as scheduled, and no anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 5 Successful : 5 Scheduled Reacquisitions : 7 Successful : 7 2.2 FHST Updates: Scheduled : 10 Successful : 10 2.3 Operations Notes: The Battery #4 Capacity Test continues. Discharge terminated nominally at 15 volts by the special hardware test limit at 175/1319Z. Recharge began at 175/1412Z (normal orbit day entry). Charge acceptance behavior is typical relative to that seen in other batteries. Capacity recovery is nominal. Using ROP SR-1, the SSR EDAC error counter was cleared three times. A real-time map was commanded at 175/1400Z. The NSSC-1 status buffer was dumped and reset at 175/1622Z per ROP NS-2. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH